OpenStudy (anonymous):

The area of a trapezoid is 200 square inches. The height of the trap is 8 inches. One of the bases is 12 inches long. Find the length of the oth base??? i got 6.5 inches but someone else told me it was 38 inches...

OpenStudy (anonymous):

so assume the smaller base is 12inches long. height is 8. split the trapezoid into two triangles and a square the square would measure 8*12 = 96. 200-96 = 104, now you would divide 104 by two to get 52 to show what each triangle must equal. Now divide that by 8 to get how much you will need on the bottom base of the trapezoid. Now the figure it going to look like the one attached in my next post, but the other base would equal 12+(6.5*2) = 25.

OpenStudy (anonymous):